HAPPINESS:

A chemistry teacher needs to mix a 20% salt solution with a 80% salt solution to make 25 qt of a 44% salt solution. How many quarts of each solution should the teacher mix to get the desired result?

Mercury:

set up a system of equations let x = number of qts from the 20% salt solution let y = number of qts from the 80% salt solution one equation will describe the number of total qrts and the other will describe the total amount of salt

Mercury:

x and y must add up to 25, therefore x + y = 25 the total amount of salt is 44% times 25, or (0.44*25). each individual solution's contribution to the mixture can be described as (percent salt content)*amount of solution therefore 0.80x + 0.20y = 0.44*25

Mercury:

so in the end you must solve x + y = 25 0.80x + 0.20y = 0.44*25 by elimination, substitution, or graphing